OpenStudy (anonymous):

I need help ;-; Which numbers are greater than 7 × 10−4? Choose all answers that are correct. A. 8 × 10−4 B. 9 × 10−5 C. 4 × 10−4 D. 3 × 10−2 E. 7 × 10−5

OpenStudy (texaschic101):

7 x 10 - 4 = 70 - 4 = 66 did you mean that, or did you mean 7 x 10^-4 ??

OpenStudy (anonymous):

the answer is a=76 and b=85

OpenStudy (anonymous):

those are the ones that are greater

OpenStudy (anonymous):

texas i ment 7 x 10^-4 :P

OpenStudy (anonymous):

ok hang on

OpenStudy (anonymous):

i needed to know that *facepalm*

OpenStudy (anonymous):

texas ill let u have it i gtg

OpenStudy (texaschic101):

7 x 10 ^ -4 = 7 x 1/10^4 = .0007 8 x 10^-4 = 8 x 1/10^4 = .0008 9 x 10^-5 = 9 x 1/10^5 = .00009 4 x 10^-4 = 4 x 1/10^4 = .0004 3 x 10^-2 = 3 x 1/10^2 = .03 7 x 10^-5 = 7 x 1/10^5 = .00007 greater ones are : A and D

OpenStudy (anonymous):

thank you so much. i was about to get this wrong :)